ABC is an equilateral triangle of side 3 units. The points P, Q lie on BC, CA re￾spectively and are such that AQ = CP = 2units. If the point R lies on AB produced

so that BR = 1unit, prove that P, Q, R are collinear


1
Expert's answer
2021-02-02T04:24:12-0500

"AR = AB+BR = 3+1 =4"

"\text{by the cosine theorem:}"

"QR^2 = AR^2+AQ^2-2AR*AQ*\cos{\angle{A}}"

"QR^2 = 4^2+2^2-2*4*2*\cos{60^0}=12"

"QR =\sqrt{12}=2\sqrt{3}"

"\text{by the cosine theorem:}"

"QP^2 = CP^2+CQ^2-2CP*CQ*\cos{\angle{C}}"

"CQ=AC-AQ=3-2=1"

"QP^2 = 2^2+1^2-2*2*1*\cos{60^0}=3"

"QP=\sqrt {3}"

"\text{by the cosine theorem:}"

"PR^2 = BP^2+BR^2-2BP*BR*\cos{\angle{(RBP)}}"

"\angle{(RBP)}= 180^0-\angle{B}= 120^0"

"BP = BC-CP = 3-2 =1"

"PR^2=1^2+1^2-2*1*1*\cos120^0=3"

"PR= \sqrt{3}"

"QR = PR+PQ"

"\text{hence P lies strictly between R,Q}"

"\text{proven}"

Answer:"\text{proven}"